Solve. 4(7x + 10) = 3(9x – 5) + 20
Answer by bubblz190(11)

First, distribute.
28x + 40 = 27x - 15 + 20
Then combine like terms.
28x + 40 = 27x + 5
x + 40 = 5
x = -35